5. What must be tge charge of an isolated spherical conductor 30.0 cm in diameter to produce an electric field of 150 N/C just outside the surface?


1
Expert's answer
2022-03-31T11:49:00-0400
"E=k\frac{q}{r^2}"

"q=\frac{Er^2}{k}=\frac{150*0.15^2}{9*10^9}=3.75*10^{-10}\:\rm C"